OpenStudy (anonymous):

HELPP ME!!WILL AWARD ...What is the equation for a circle with center (2, -3) and radius of 6?

OpenStudy (anonymous):

\[r^2 = (x - h)^2 + (y-k)^2\]

OpenStudy (tanya123):

(x - h)^2 + (y - k)^2 = r^2 where (h,k) is the center and r is the radius Substitute the center (h,k)=(0,2) and r = SR(6) (x - 0)^2 + (y - 2)^2 = ((SR(6))^2 x^2 + y^2 - 4y + 4 = 6 x^2 + y^2 - 4y = 2

OpenStudy (anonymous):

(x - 2)2 + (y + 3)2 = 6 (x - 2)2 + (y + 3)2 = 36 (x + 2)2 + (y + 3)2 = 36 (x - 2)2 + (y - 3)2 = 36

OpenStudy (anonymous):

@tanya123 @nydr ^ that is the answer choices.

OpenStudy (anonymous):

good to know:)

OpenStudy (anonymous):

@nydr so which one is it?

OpenStudy (anonymous):

(x-2)^2 + (y+3)^2 = 36

OpenStudy (anonymous):

is the real one

OpenStudy (anonymous):

so it would be x^2 - 4x +4 + y^2 + 6y + 9 =36 x^2 +y^2 -4x +6y +13-36 x^2 + y^2 - 4x +6y - 23 = 0